# Deep-learning-based continuous attacks on QKD protocols

### v3

@@ Abstract
The most important characteristic of a Quantum Key Distribution (QKD) protocol is its security against third-party attacks, and the potential countermeasures available. While new types of attacks are regularly developed in the literature, they rarely involve the use of weak continuous measurement and more specifically machine learning to infer the qubit states. In this paper, we design a new individual attack scheme called Deep-learning-based continuous attack (DLCA) that exploits continuous measurement together with the powerful pattern recognition capacities of deep recurrent neural networks. We show that, when applied to the BB84 protocol, our attack increases only slightly the Quantum Bit Error Rate (QBER) of a noisy channel and allows the spy to infer a significant part of the sifted key. In addition, we show it yields similar performances in terms of information gain when compared to an optimal individual attack, namely the phase-covariant quantum cloner. Our individual attack scheme demonstrates deep-learning-enhanced quantum state tomography applied to QKD and could be generalized in many different ways,
@@ I Introduction
In the last decades, the demand for fast, secure, and reliable data connections has significantly increased. To meet this demand, it is essential to enhance the computational power of network systems through high-performance technologies. Quantum computing is one such technology, showing a clear potential to outperform current classical computing systems. Quantum computing-assisted communications have therefore been extensively studied and developed in recent years, and hold great promise for improving communications and security in today’s networks [1, 2].
@@
where $N_{\mathrm{total}}$ is the total number of qubits received where Bob used the right measurement basis, and $N_{\mathrm{error}}$ is the number of incorrect results he gets among these qubits. In the case of a perfect quantum communication channel, the QBER should be zero. However, in the presence of a spy, the qubits states are usually altered and the QBER non-zero despite Alice and Bob using the same basis for the qubits, which should thus signal the two parties that something went wrong. Complications then arise because in practice, the quantum channel is not perfectly isolated from its environment (e.g., the optical fiber could be leaky [26, 27, 28]) and Bob measurement apparatus could be defective, contributing to another cause of QBER enhancement. Hence, distinguishing an attack from intrinsic noise in the channel is not always easy. Despite this, several fundamental papers established the unconditional security of the BB84 protocol in the early 2000s [29, 30, 5, 31, 32]. The security is unconditional in the sense that no assumption is made on Eve’s attacks: Eve can perform any measurement scheme on the channel and the channel may be subject to dissipation, Alice and Bob will upper-bound Eve’s obtainable information and, provided the QBER is below a certain threshold, reduce it to an arbitrary low level using privacy amplication.
@@
However, only a couple of works have investigated how artificial intelligence could be used to develop more effective attacks: a convolutional neural network was trained to help the eavesdropper choose the best opportunity to launch an entanglement-distillation-attack [43], an ML algorithm was shown to be able to analyze the power originating from the integrated electrical control circuit to perform a power-analysis-attack [44], and a quantum circuit implementation of the BB84 protocol was interpreted as a quantum machine learning task, allowing to find a cloning algorithm outperforming known ones [45]. Also, a deep convolutional neural network was used to monitor the electromagnetic emissions of a QKD emitter (Deep-learning-based radio-frequency side-channel attack [46]). Despite neural networks demonstrating a better ability to perform quantum state reconstruction using partial information and fewer measurements than classical state tomography [47, 48, 49, 50], the literature does not show extensive research on ML/DL-based attacks.
@@
In this paper, we develop a new type of individual attack based on continuous measurement [51] of single polarized photons and apply it for concreteness in the context of the BB84 protocol, by contrast to other works that usually apply this kind of measurement on CVQKD. The general motivation is to evaluate the performance of attacks that produce only a small perturbation to the qubits, by contrast with the effects of projective measurement usually involved in the other types of attacks. In particular, we investigate how the effects of the spy measurement can be optimally hidden by the intrinsic noise of the quantum communication channel by minimizing the increase in QBER due to the measurement. At the same time, we investigate how neural networks can effectively use the information extracted by these types of measurement, to infer a more significant part of the key than conventional means. To do so, we feed the outcome of the continuous measurement, also called homodyne photo currents, to a Long Short-Term Memory (LSTM) recurrent neural network [52] to retrieve the initial states of the photons sent by Alice, which compose the sifted key generated. The main point of this paper is not to question the security of the BB84 protocol, but rather to show an application of deep learning in QKD, and to raise awareness on its use in the context of attacks.
@@
This paper is organized as follows: In Sec. II we first summarize the BB84 protocol, present our model of the qubit dynamics in the quantum communication channel when subject to intrinsic dissipation and continuous measurement, investigate how a spy could use the outcome of this measurement to obtain the initial state of the qubit and present the neural network we implemented to do so. In Sec. III we compare the results obtained via a basic projective measurement (Intercept-and-Resend attack) and our measurement scheme. We then extend the theory to include quantum feedback, to try to decrease the impact of the measurement while keeping the same accuracy of the neural network. In Sec. IV., we discuss our attack scheme in terms of information gain and how it fits into the thresholds established by the information-disturbance principle, and calculate the typical key rate Alice and Bob should achieve to secure the protocol. In particular, we compare our attack performances against the ones of an optimal individual attack strategy for the BB84 protocol: the covariant-phase quantum cloner [53, 54]. Finally, in Sec. V, we conclude and discuss potential perspectives of our work.
@@ II.1 BB84 protocol
The BB84 protocol, sketched in Fig. 1, implements a shared secret key between two parties by storing private key bits in linearly polarized states of photons. There are four initial states: vertically and horizontally polarized states represented by |$\ket{0}$ and |$\ket{1}$ respectively, and two diagonally polarized states defined as
$\ket{+}$ =|$=\frac{\ket{0}+\ket{1}}{\sqrt{2}},$ (2)
$\ket{-}$ =|$=\frac{\ket{0}-\ket{1}}{\sqrt{2}}.$
@@ 1. Step 1:
    Alice chooses a random data bit string $b$ (e.g., $b=01011\dots)$. She encodes each data bit randomly as the quantum states |$\ket{0}$ or $\ket{+}$ if the corresponding bit of $b$ is 1 and |$\ket{1}$ or $\ket{-}$ if the corresponding bit of $b$ is 0.
@@ II.2 Dissipative qubit dynamics conditioned on measurement
Figure 1: Sketch representation of our attack scheme applied on the BB84 protocol. Alice sends linearly polarized photons to Bob via an optical fiber, while traveling they are subject to intrinsic dissipation in the fiber with a rate $\gamma_{D}$ and to weak measurement with a rate $\gamma_{E}$ performed by Eve on a certain portion of the fiber, the output of which being treated by a neural network to determine the initial state. In the sketched example, Alice sends six photons to Bob, among which three are polarized in the Pauli-X eigenbasis and three in the Pauli-Z eigenbasis. Bob measures randomly in one of these two bases each of the photons received, and the basis chosen is the right one for four of these. However, even if the measurement basis is right, they are two errors. The sketch highlights the two possible mechanisms for errors: intrinsic dissipation (blue dashed square) and measurement by a third party (purple dashed square). In this example, if Alice and Bob compare all the measurements when they chose the same basis, the QBER [Eq. (1)] would be $50\%$.
For concreteness, we model the dynamics of each individual qubit in the quantum communication channel as subjected to i) intrinsic dissipation acting on the channel and ii) a continuous measurement performed by a spy (see Fig. 1). More specifically, we consider the following stochastic master equation (written here in Itô form)
@@
where $H$ is the channel Hamiltonian defined as $H=\omega\sigma_{z}$ for the initial states |$\ket{0}$ or |$\ket{1}$ and $H=\omega\sigma_{x}$ for $\ket{+}$ or $\ket{-}$ [26] with $\sigma_{x}=|0\rangle\langle 1|+|1\rangle\langle 0|$ and $\sigma_{z}=|0\rangle\langle 0|-|1\rangle\langle 1|$ the standard Pauli operators, where $\rho_{J}$ is the density operator of the qubit conditioned on the measurement with efficiency $\eta\in[0,1]$ of the homodyne current [60]
$Jdt=\sqrt{\gamma_{E}}\langle e+e^{\dagger}\rangle dt+\frac{dW}{\sqrt{\eta}},$ (4)
@@
for a given operator $o$. In Eq. (3), the first line represents the effect of the unitary dynamics of the channel governed by the Hamiltonian $H$ as well as the effect of the intrinsic dissipation produced by the operator $d$ occuring at rate $\gamma_{D}$, while the second line represents the effect of the eavesdropping produced by the operator $e$ at rate $\gamma_{e}$, which decomposes into an incoherent term and a non-linear stochastic term, where $dW$ is a Wiener increment satisfying $dW^{2}=dt$. For concreteness, we set throughout this work the dissipation operator to be
@@ II.4 Neural network quantum tomography based on the measurement
Figure 2: Sketch representation of the LSTM architecture used in this paper. The input layer is composed of $40$ LSTM units, and is followed by two dense hidden layers of $40$ ReLU neurons and $20$ sigmoid neurons respectively. The output layer is composed of $4$ softmax neurons. On the left is a representation of one LSTM recurrent unit, composed of three gates with sigmoid activation functions (forget, input and output). These 3 gates determine which information from the prior hidden state must be erased, taken into account and stored respectively.
The input layer of our network is a LSTM one with $40$ units in its hidden state, to take as input the time series data that are the photo currents. We use the last hidden state (dimension $40$) as the input of a $40$-neurons dense layer, followed by another $20$-neurons dense layer, with activation functions set to ReLU and sigmoid respectively. The output layer is a $4$ neurons softmax layer. The loss function we use for training is the sparse categorical cross entropy since we deal with a $4$-class classification problem, and the optimizer is Adam with default parameters. The training consists in one epoch over $9\times 10^{4}$ photo currents while the testing is done on $10^{4}$ photo currents. Before entering the network the currents are standardized and flipped along their time axis.
@@ II.5 Impact of measurement VS Accuracy
As explained earlier, the goal of the spy is to minimize their impact on the photon states while maximizing the part of the sifted key obtained. To quantify the impact of the measurement, we use the QBER introduced earlier [Eq. (1)], as it is directly measurable by Alice and Bob and allows them to assess the security of the protocol. To quantify the success of the spy in retrieving the initial state of the photons, we use the spy accuracy that we will denote by $A$. In the context of our deep learning approach, this is the neural network test accuracy [72] which is defined as the percentage of good predictions among all the predictions of the network on the test set. However, when dealing with a projective measurement, this accuracy is defined as the probability that the spy measures the right state. Note that the amount of information extracted from the qubits (e.g., the information gain) is defined by the measurement scheme. On the other hand, accuracy $A$ depends on both the information gain and the ability of the neural network to efficiently harness it, as detailed in Sec. IV.1. Thus $A$ does not correspond properly to a measure of the extracted information in the sense of Shannon [73]. However, we use it for convenience to quantify the success of the eavesdropping scheme, since it is a performance measure that represent the average percentage of the sifted key obtained by the spy.
@@ III Results
In this section, we study the impact of our attack and its performance in different cases in terms of QBER and accuracy $A$. We first compute the QBER in the case of no attack. Then, we study a simple standard projective measurement attack, before investigating our continuous measurement scheme. We also study the performance of quantum feedback the spy could use to cover their tracks. Note that in this section, except stated otherwise, all time durations are measured in units of $1/\gamma_{D}$.
@@ III.1 No attack
where $t_{f}$ is the total travel time of the qubit in the noisy quantum channel. Hence, the QBER ranges from 00 for a perfect channel to $25\%$ for a very noisy channel or a very long travel time. This result is displayed as the black dashed line in Fig 3.
@@ III.2 Attack via projective measurement
In this case, the accuracy $A$, can be calculated exactly by solving the Lindblad master equation with a single jump operator $L=\sigma_{x}$ (see Appendix C) and reads
@@
Interestingly, we see that the time $t^{*}$ at which Eve performs her measurement does not impact the probability that Bob measures the state he is supposed to. Also, comparing Eqs. (8) and (10), we clearly see that Alice and Bob will easily distinguish the presence of the spy from intrinsic dissipation. This result is displayed as the dotted black line in Fig 3.
Figure 3: Top: QBER as a function of $\gamma_{D}t$ for different attack schemes and angles $\theta$. The black dashed line represents the QBER when no measurement is performed (i.e., produced only by intrinsic dissipation) [Eq. (8)]. The dotted black line is the QBER with dissipation and projective measurement [Eq. (10)] and the red line is the QBER caused by a continuous measurement with $e=\sigma_{z}$ ($\theta=\pi/2$). The orange line is the QBER when using the angle $\theta_{A}^{*}$ maximizing the accuracy (which we update every $\gamma_{D}t=0.3$), the green line for the angle $\theta_{\mathrm{qber}}^{*}$ minimizing the QBER, and the blue line for the angle $\theta_{\lambda}^{*}$ minimizing $\lambda$. Bottom: optimized angles with respect to the QBER itself (green line), the NN test accuracy (orange line) and $\lambda$ (blue line).
@@ III.3 Attack via continuous measurement
and consider in the first instance that the homodyne measurement is performed during the whole travel time, set to $\gamma_{D}t_{f}=3$, and with other parameters $\eta=0.5$, $\omega=\gamma_{E}=\gamma_{D}$. With these parameters we obtain, from the solutions of Eq. (3), a QBER of $49\%$ (red line in Fig 3), much higher than the $37\%$ of the projective measurement obtained from Eq. (10) and the $24.4\%$ of the case with dissipation only, obtained from Eq. (8). In addition, we get a neural network test accuracy $A\approx 71\%$, which is within the interval given by Eq. (9). We interpret this result as coming from the fact that the states |$\ket{0}$ and |$\ket{1}$ being eigenstates of the measurement operator $e=\sigma_{z}$, the neural network can approximately distinguish these two while it must do random guess on the other two states ($\ket{+}$ and $\ket{-}$), which would lead to a test accuracy around $75\%$ (see Appendix E). This suggests that the neural network puts a heavy weight on the early-time behavior of the photo current, as confirmed in Appendix F. Hence, we see that the spying accuracy achieved via this simple continuous measurement scheme is comparable to the one achieved via the projective measurement, but the QBER is higher.
@@
Let us first look at the accuracy $A$ yielded by this new measurement operator as a function of $\theta/\pi$, which is depicted in Fig. 4, together with the associated standard deviation. There are four angles leading to an accuracy around $90\%$, as summarized in Table 1, which is much higher than the $71\%$ found earlier. Note that the four angles seem equivalent given their values and the standard deviations.
Figure 4: Mean estimated accuracy $A$ (blue) and standard deviation (orange) of the model on the test set as a function of $\theta$. The photo currents of the test set were obtained using Eq. (3) and (4). Circled in red are the four maximum accuracy values and their corresponding standard deviations, reported in Table I. Other parameters are $\gamma_{D}t_{f}=3$ and $\omega=\gamma_{E}=\gamma_{D}$.
@@
To obtain the impact of the measurement on the BB84 protocol itself, we average the QBER over the four possible initial states, and evaluate it as a function of $\theta$ and $\gamma_{D}t$, as displayed in Fig. 5.
Figure 5: QBER as a function of time and measurement angle $\theta$. The evolution of the photon states through time was obtained using Eq. (3). Other parameters are $\gamma_{D}t_{f}=3$ and $\omega=\gamma_{E}=\gamma_{D}$.
One can see there is a trade-off between the accuracy and the disturbance that the measurement induces. Indeed, the angles that minimize the QBER, which are $\theta=0\pm k\pi,k\in\mathbb{Z}$, yield a low test accuracy, thus decreasing the spy accuracy (see Fig. 4). In order to quantify this tradeoff, we define a new quantity $\lambda(\theta)$ as the QBER divided by the accuracy $A$ of the network for a given measurement basis (i.e., a given $\theta$)
@@
which is shown for $\gamma_{D}t_{f}=3$ in Fig. 6. We see that among the four measurement angles maximizing the spy accuracy (circled in red), $\theta=1.86\pi$ yields the lowest $\lambda$ ratio, with an accuracy around $90\%$ and a QBER around $37\%$, though the other angles give similar performances.
@@
Figure 6: $\lambda(\theta)$ [Eq. (13)] as a function of $\theta$. Circled in red are the four values of $\theta$ maximizing $A$ found in Fig. 4 and reported in Table I. Other parameters are $\gamma_{D}t_{f}=3$ and $\omega=\gamma_{E}=\gamma_{D}$.
The top panel of Fig. 3 gathers all the QBER discussed above. In addition, three other lines are present: the QBER for a continuous measurement when $\theta$ is set to minimize the QBER itself (green line), to maximize the accuracy (orange line), and finally to minimize $\lambda$ (blue line). In the bottom panel, the $\theta$ values of these three lines are represented, along with the four optimal angles from Table 1 as guides to the eye (black dashed lines). The QBER optimized with respect to itself is at first higher than the one from no measurement at all, before converging to the same value of $25\%$. On the lower panel, we see the angle $\theta^{*}_{qber}$ is always equal to 00. This is due to the fact that it corresponds to $e=\sigma_{x}$, which would act in a similar way as the intrinsic dissipation, and thus make the state converge faster to the same steady state yielding a QBER of $25\%$, which is the same value as with no measurement, thus the minimal value a measurement can yield. However, even though Alice and Bob could not in average detect this eavesdropping (at least for large $\gamma_{D}t_{f}$), the neural network manages to learn something from it since its mean accuracy for this case ($\theta=0$, i.e., $e=\sigma_{x}$) is higher than $25\%$ (see Fig. 4). Also, the measurement optimized with respect to the accuracy yields a lower QBER than a projective measurement for $\gamma_{D}t_{f}\leq 2.5$ and the one optimized with respect to $\lambda$ yields an even lower QBER, but still higher than the green line.
Finally, to make the measurement more realistic, one can reduce the duration of the continuous measurement so that it starts later than the initial time and finishes before the final time. We choose here the optimal measurement angle found earlier, $\theta=1.86\pi$, and set it to start at $\gamma_{D}t=0.1$ and define a duration $\Delta t$ of the continuous measurement with the aim of decreasing its impact on the qubit states while maintaining a reasonable accuracy. As shown in Fig. 7, which displays $A$ as a function of $\gamma_{D}\Delta t$, the accuracy reaches $86.1\%$ with a reasonable standard deviation ($2\%$) when measuring between $\gamma_{D}t=0.1$ and $0.5$. For longer duration, we see that the accuracy is reasonable but the standard deviation becomes very high as compared to Fig. 4. This suggests that the initial time-evolution (i.e., before $\gamma_{D}t=0.1$), which is not fed to the neural network here, is crucial to make it sharp. Altogether, taking $\theta=1.86\pi$ and the optimal duration $\gamma_{D}\Delta t=0.4$, we obtain
$A=86.1\%,$ (14)
$\mathrm{QBER}=27.6\%,$ (15)
the latter representing only a $2.6\%$ increase compared to the time-evolved state where no measurement is made. Comparatively, the Time shift [16, 17] attack scheme yields an accuracy around $60-70\%$ and an increase in QBER between $1$ and $2\%$. Hence, though the noise model for the fiber is not the same, our new attack scheme seems to provide similar QBER increases and very high accuracies. A summary of the results of this section can be found in Table 2.

### v4

@@ Abstract
The most important characteristic of a Quantum Key Distribution (QKD) protocol is its security against third-party attacks, and the potential countermeasures available. While new types of attacks are regularly developed in the literature, they rarely involve the use of weak continuous measurement and more specifically machine learning to infer the qubit states. In this paper, we design a new individual attack scheme called Deep-learning-based continuous attack (DLCA) that exploits continuous measurement together with the powerful pattern recognition capacities of deep recurrent neural networks. As a minimal model, we present its performances when applied in the case of the BB84 protocol with intrinsic noise in the communication channel. Our results suggest that our attack’s performances lie between the ones of standard intercept-and-resend attacks and of the optimal individual attack, namely the phase-covariant quantum cloner. Our attack scheme demonstrates deep-learning-enhanced quantum state tomography applied to QKD, and could be generalized in many different ways, notably in the cases of quantum hacking attacks targeting implementation vulnerabilities that could compromise the security of QKD protocols.
@@ I Introduction
In the last decades, the demand for fast, secure, and reliable data connections has significantly increased. To meet this demand, it is essential to enhance the computational power of network systems through high-performance technologies. Quantum computing is one such technology, showing a potential to outperform current classical computing systems. Quantum computing-assisted communications have therefore been extensively studied and developed in recent years, and hold great promise for improving communications and security in today’s networks [1, 2].
@@
where $N_{\mathrm{total}}$ is the total number of qubits received where Bob used the right measurement basis, and $N_{\mathrm{error}}$ is the number of incorrect results he gets among these qubits. In the case of a perfect quantum communication channel, the QBER should be zero. However, in the presence of a spy, the qubits states are usually altered and the QBER non-zero despite Alice and Bob using the same basis for the qubits, which should thus signal the two parties that something went wrong. Complications then arise because in practice, the quantum channel is not perfectly isolated from its environment (e.g., the optical fiber could be leaky [26, 27, 28]) and Bob measurement apparatus could be defective, contributing to another cause of QBER enhancement. Hence, distinguishing an attack from intrinsic noise in the channel is not always easy. Despite this, several fundamental papers established the unconditional security of the BB84 protocol in the early 2000s [29, 30, 5, 31, 32]. The security is unconditional in the sense that no assumption is made on Eve’s attacks: Eve can perform any measurement scheme on the channel and the channel may be subject to dissipation, Alice and Bob will upper-bound Eve’s obtainable information and, provided the QBER is below a certain threshold, reduce it to an arbitrarily low level using privacy amplification.
@@
However, only a couple of works have investigated how artificial intelligence could be used to develop more effective attacks: a convolutional neural network was trained to help the eavesdropper choose the best opportunity to launch an entanglement-distillation-attack [43], an ML algorithm was shown to be able to analyze the power originating from the integrated electrical control circuit to perform a power-analysis-attack [44], and a quantum circuit implementation of the BB84 protocol was interpreted as a quantum machine learning task, allowing to find a cloning algorithm outperforming known ones [45]. Also, a deep convolutional neural network was used to monitor the electromagnetic emissions of a QKD emitter (Deep-learning-based radio-frequency side-channel attack [46]). Despite neural networks demonstrating a better ability to perform quantum state reconstruction using partial information and fewer measurements than classical state tomography [47, 48, 49, 50], the literature does not show extensive research on ML/DL-based attacks. Often, Eve is modeled as introducing an ancillary quantum system (i.e., a probe) that interacts unitarily with the traveling qubits through some channel, before either being measured directly (individual attacks) or stored in order to perform a coherent or collective measurement later. The power of the attack is evaluated by an information-theoretic upper bound, such as the mutual information between the probe system and Alice’s bits which represents the maximum information Eve could, in principle, extract, regardless of the specific measurement.
@@
In this paper, we develop a new type of individual attack based on continuous measurement [51] of single polarized photons and apply it for concreteness in the context of the BB84 protocol, by contrast to other works that usually apply this kind of measurement on CVQKD. The general motivation is to evaluate the performance of a specific measurement scheme that theoretically produces only a small perturbation to the qubits, by contrast with the effects of projective measurement usually involved in the other types of attacks. In particular, we investigate how the effects of the spy measurement can be optimally hidden by the intrinsic noise of the quantum communication channel by minimizing the increase in QBER due to the measurement. At the same time, we investigate how neural networks can effectively use the information extracted by these types of measurement, to infer a more significant part of the key than conventional means. To do so, we feed the outcome of the continuous measurement, also called homodyne photo currents, to a Long Short-Term Memory (LSTM) recurrent neural network [52] to retrieve the initial states of the photons sent by Alice, which compose the sifted key generated. The main point of this paper is not to question the security of the BB84 protocol, but rather to show an application of deep learning in QKD, namely deep-learning-assisted quantum state tomography, and to raise awareness on its potential use in the context of attacks.
@@
This paper is organized as follows: In Sec. II we first summarize the BB84 protocol, present our model of the qubit dynamics in the quantum communication channel when subject to intrinsic dissipation and continuous measurement, investigate how a spy could use the outcome of this measurement to obtain the initial state of the qubit and present the neural network we implemented to do so. In Sec. III we compare the results obtained via a basic projective measurement (Intercept-and-Resend attack) and our measurement scheme. In Sec. IV., we discuss our attack scheme in terms of information gain and how it fits into the thresholds established by the information-disturbance principle, and calculate the typical key rate Alice and Bob should achieve to secure the protocol. In particular, we compare our attack performances against the ones of an optimal individual attack strategy for the BB84 protocol: the covariant-phase quantum cloner [53, 54]. Finally, in Sec. V, we conclude and discuss potential perspectives of our work.
@@ II.1 BB84 protocol
[…]
@@ 1. Step 1:
[…]
@@ II.2 Dissipative qubit dynamics conditioned on measurement
Figure 1: Sketch representation of our attack scheme applied on the BB84 protocol. Alice sends linearly polarized photons to Bob via an optical fiber, while traveling they are subject to intrinsic dissipation in the fiber with a rate $\gamma_{D}$ and to weak measurement with a rate $\gamma_{E}$ performed by Eve on a certain portion of the fiber, the output of which being treated by a neural network to determine the initial state. In the sketched example, Alice sends six photons to Bob, among which three are polarized in the Pauli-X eigenbasis and three in the Pauli-Z eigenbasis. Bob measures randomly in one of these two bases each of the photons received, and the basis chosen is the right one for four of these. However, even if the measurement basis is right, there are two errors. The sketch highlights the two possible mechanisms for errors: intrinsic dissipation (blue dashed square) and measurement by a third party (purple dashed square). In this example, if Alice and Bob compare all the measurements when they chose the same basis, the QBER [Eq. (1)] would be $50\%$.
For concreteness, we model the dynamics of each individual qubit in the quantum communication channel as subjected to i) intrinsic dissipation acting on the channel and ii) a continuous measurement performed by a spy (see Fig. 1). More specifically, we consider the following stochastic master equation (written here in Itô form) [60]
@@
[…]
$Jdt=\sqrt{\eta\gamma_{E}}\langle e+e^{\dagger}\rangle dt+dW,$ (4)
@@
for a given operator $o$. In Eq. (3), the first line represents the effect of the unitary dynamics of the channel governed by the Hamiltonian $H$ as well as the effect of the intrinsic dissipation produced by the operator $d$ occuring at rate $\gamma_{D}$, while the second line represents the effect of the eavesdropping produced by the operator $e$ at rate $\gamma_{e}$, which decomposes into an incoherent term and a non-linear stochastic term, where $dW$ is a Wiener increment satisfying E[dW]=0[dW]=0 and $dW^{2}=dt$. For concreteness, we set throughout this work the dissipation operator to be
@@ II.4 Neural network quantum tomography based on the measurement
Figure 2: Sketch representation of the LSTM architecture used in this paper. The input layer is composed of $100$ LSTM units, and is followed by two dense hidden layers of $40$ ReLU neurons and 44 linear neurons respectively. On the left is a representation of one LSTM recurrent unit, composed of three gates with sigmoid activation functions (forget, input and output). These 3 gates determine which information from the prior hidden state must be erased, taken into account and stored respectively.
The input layer of our network is a LSTM one with $100$ units in its hidden state, to take as input the time series data that are the photo currents. We use the last hidden state along the sequence length (dimension $100$) as the input of a $40$-neuron dense layer with activation function set to ReLU. The output layer is a $4$-neuron linear layer. The loss function we use for training is the sparse categorical cross entropy since we deal with a $4$-class classification problem, and the optimizer is Adam with default parameters. The model is trained on $9\times 10^{4}$ photo currents, with dropout to reduce overfitting, and tested on $10^{4}$ photo currents. Before being fed into the network, the photocurrent values at each time step are standardized to have zero mean and unit variance.
@@ II.5 Impact of measurement VS Accuracy
As explained earlier, the goal of the spy is to minimize their impact on the photon states while maximizing the part of the sifted key obtained. To quantify the impact of the measurement, we use the QBER introduced earlier [Eq. (1)], as it is directly measurable by Alice and Bob and allows them to assess the security of the protocol. To quantify the success of the spy in retrieving the initial state of the photons, we use the spy accuracy that we will denote by $A$. In the context of our deep learning approach, this is the neural network test accuracy [72] which is defined as the percentage of good predictions among all the predictions of the network on the test set. Although there are several metrics used in machine learning (e.g., F1-score), accuracy reflects the model percentage of success in a given task, and suits our problem given the four initial states are equiprobable.
When dealing with a projective measurement, this accuracy is defined as the probability that the spy measures the right state. Note that the amount of information extracted from the qubits (e.g., the information gain) is defined by the measurement scheme. On the other hand, accuracy $A$ depends on both the information gain and the ability of the neural network to efficiently harness it, as detailed in Sec. IV.1. Thus $A$ does not correspond properly to a measure of the extracted information in the sense of Shannon [73]. However, we use it for convenience to quantify the success of the eavesdropping scheme, since it is a performance measure that represents the average percentage of the sifted key obtained by the spy. The accuracy as defined is a performance measure for the $4$-states classification task, and does not entirely reflect the real performance of the spy, which is to obtain the sifted key, i.e., a $2$-class classification problem. For this purpose, the key accuracy $A_{key}$ is employed here.
@@ III Results
In this section, we study the impact of our attack and its performance in different cases in terms of QBER and accuracy $A$. We first compute the QBER in the case of no attack. Then, we study a simple standard projective measurement attack, before investigating our continuous measurement scheme. Note that in this section, except stated otherwise, all time durations are measured in units of $1/\omega$.
@@ III.1 No attack
where $t_{f}$ is the total travel time of the qubit in the noisy quantum channel. Hence, the QBER ranges from 0 for a perfect channel to $25\%$ for a very noisy channel or a very long travel time
@@ III.2 Attack via projective measurement
In this case, the accuracy $A$ can be calculated exactly by solving the Lindblad master equation with a single jump operator $L=\sigma_{x}$ (see Appendix C) and reads
@@
Interestingly, we see that the time $t^{*}$ at which Eve performs her measurement does not impact the probability that Bob measures the state he is supposed to. Also, comparing Eqs. (8) and (10), we clearly see that Alice and Bob will easily distinguish the presence of the spy from intrinsic dissipation.
@@ III.3 Attack via continuous measurement
and consider in the first instance that the homodyne measurement is performed during the whole travel time, set to $\omega t_{f}=0.1$, and with other parameters $\eta=1$, $\gamma_{E}=5\gamma_{D}=5\omega$. With these parameters we obtain, from the solutions of Eq. (3), a QBER of $20.5\%$, lower than the $37\%$ of the projective measurement obtained from Eq. (10) and higher than the $4.5\%$ of the case with dissipation only, obtained from Eq. (8). In addition, we get a neural network test accuracy $A\approx 43\%$, which is below the interval given by Eq. (9). Hence, we see that the spying accuracy achieved via this simple continuous measurement scheme is lower than the one achieved via the projective measurement, but the QBER is lower.
@@
Let us first look at the accuracy $A$ yielded by this new measurement operator as a function of $\theta/\pi$, which is depicted in Fig. 3, together with the associated standard deviation. There are four angles leading to an accuracy around $46\%$, as summarized in Table 1, which is higher than the $43\%$ found earlier. Note that the four angles seem equivalent given their values and the standard deviations.
Figure 3: Mean estimated accuracy AA (blue) and standard deviation (orange) of the model on the test set as a function of $\theta$. The photo currents of the test set were obtained using Eq. (3) and (4). Circled in red are the four maximum accuracy values and their corresponding standard deviations, reported in Table I. Other parameters are $\omega t_{f}=0.1$ and $\gamma_{E}=5\omega=5\gamma_{D}$.
@@
To obtain the impact of the measurement on the BB84 protocol itself, we average the QBER over the four possible initial states, and evaluate it as a function of $\theta$ and $\omega t$, as displayed in Fig. 4.
Figure 4: QBER as a function of time and measurement angle $\theta$. The evolution of the photon states through time was obtained using Eq. (3). Other parameters are $\omega t_{f}=0.1$ and $\gamma_{E}=5\omega=5\gamma_{D}$.
One can see there is a trade-off between the accuracy and the disturbance that the measurement induces. However, the angles that minimize the QBER, which are $\theta=0\pm k\pi,k\in\mathbb{Z}$, yield a better test accuracy, thus increasing the spy accuracy (see Fig. 3). In order to quantify this tradeoff, we define a new quantity $\lambda(\theta)$ as the QBER divided by the accuracy $A$ of the network for a given measurement basis (i.e., a given $\theta$)
@@
which is shown for $\omega t_{f}=0.1$ in Fig. 5. As expected, we observe that among the four measurement angles maximizing the spy accuracy (circled in red), $\theta=\pi$ yields the lowest $\lambda$ ratio, with an accuracy around $46.5\%$ and a QBER around $17.5\%$, though the other angles give similar performances.
@@
Figure 5: $\lambda(\theta)$ [Eq. (13)] as a function of $\theta$. Circled in red are the four values of $\theta$ maximizing AA found in Fig. 3 and reported in Table I. Other parameters are $\omega t_{f}=0.1$ and $\gamma_{E}=5\omega=5\gamma_{D}$.
Finally, we analyze the impact of the measurement duration (denoted $\omega\Delta t$) on the performance of the attack. Indeed, one could expect the information about the initial state to be mostly contained in the early stages of the currents, thus allowing to decrease its duration and its impact on the qubit states while maintaining a reasonable accuracy. We choose here the optimal measurement angle found earlier, $\theta=\pi$. As shown in Fig. 6, which displays $A$ as a function of $\omega\Delta t$, the accuracy reaches $44\%$ by only measuring until $\omega\Delta t=0.07$ while the QBER decreases to $14.5\%$.
Altogether, taking $\theta=\pi$, the optimal duration $\omega\Delta t=0.07$, and increasing $\gamma_{E}$ to $10\omega$ we obtain
$A=47.5\%,$ (14)
$\mathrm{QBER}=20\%,$ (15)
the latter representing a $15\%$ increase compared to the time-evolved state where no measurement is made, as summarized in Table 2. However, this accuracy over the four initial states represents an accuracy over the key bits $A_{key}=73\%$.